# Strong Gravitational Lensing Effects of the Rotating Short-Haired Black Hole and Constraints from EHT Observations

### v2

@@ V Constraints from the EHT
Figure 10: After taking the rotating short-haired BH model as a candidate model for M87* and Sgr A*, the parameter space ($Q_{m}/M,a/M$) is constrained using the EHT data. The left hand figure corresponds to M87*, and the right hand figure corresponds to Sgr A*. The red dashed line represents the boundary of the corresponding first confidence interval. Here, the inclination angle $\theta_{o}=\frac{\pi}{2}$.
@@
In the preceding section, we delved into the influence exerted by the hair parameter on the strong gravitational lensing phenomenon. The hair parameter acts by modifying the spacetime structure in the vicinity of the rotating short-haired BH, thereby inducing alterations in the propagation trajectories of light rays within the BH’s gravitational field. As light rays approach the event horizon, the intense gravitational field of the black hole causes the light rays to undergo deflection, giving rise to the gravitational lensing effect. This may potentially result in the formation of multiple images or regions of enhanced luminosity, as elaborated in detail in the previous section. With further deflection of the light rays, a black hole shadow is ultimately formed. The geometric properties of the shadow of a rotating short-haired BH are not only contingent upon the black hole’s spin but are also significantly influenced by the hair parameter. Consequently, while having an impact on the propagation paths of light rays, the hair parameter exerts a far-reaching influence on both the shape and size of the black hole shadow. A comprehensive discussion regarding the effect of the hair parameter on the shape of the shadow surrounding a rotating short-haired BH can be found in the reference Tang and Xu (2022). Therefore, our emphasis does not lie in a detailed exploration of the effects on the shadow shapes. Instead, we are committed to systematically constraining the relevant parameter space of the rotating short-haired BH by leveraging the shadow data of supermassive BHs obtained from the observations of the EHT.
@@
In 2019, the EHT successfully captured the image of the black hole shadow of the supermassive BH M87* located at the center of the elliptical galaxy M87, which brought revolutionary progress to the observational and theoretical research of supermassive BHs. The EHT team measured the diameter of the black hole shadow of M87* to be $\theta_{d}=42\pm 3\ \mu\mathrm{as}$ Akiyama et al. (2019b, a, d). This result not only provides us with precise observational evidence of the geometric structure of the BH but also offers evidence for verifying the validity of general relativity in extreme gravitational fields. Subsequently, in 2022, the EHT team further enhanced the observational technology and data processing capabilities and successfully measured the diameter of the black hole shadow of the supermassive BH Sgr A* located at the center of the Milky Way. The result is $\theta_{d}=51.8\pm 2.3\ \mu\mathrm{as}$ Akiyama et al. (2022a). These data play a crucial role in the detection of gravitational theories and parameter constraints, such as Jusufi et al. (2022); Zahid et al. (2023); Afrin et al. (2023); Raza et al. (2024); Sánchez (2024); Zahid et al. (2023).
@@
For axisymmetric black holes, the geometric shape of the black hole shadow is influenced by the spin parameter. To more accurately describe the shape of the shadow, an astronomical coordinate system Hioki and Maeda (2009) is usually introduced, which is defined as
@@
Here, the prime (’) indicates the derivative with respect to $r$, and $\Delta$ is a function term associated with the rotating short-haired BH. The detailed derivation process of this part can be found in corresponding literature, such as Tang and Xu (2022); Zahid et al. (2023); Sarikulov et al. (2022), etc.
@@
Overall, whether using the shadow data of M87* or Sgr A* to constrain the parameter space $(Q_{m}/M,a/M)$ of the rotating short-haired BH model, our analysis results show that the parameter space of the rotating short-haired BH model is highly consistent with the range of the observational data from the EHT. Therefore, through a systematic analysis of the EHT observational results of M87* and Sgr A* within the first confidence interval, we cannot rule out the rotating short-haired BH model as a potential candidate for real universe black holes. In addition, although the current observational accuracy still has limitations, the EHT data has provided a certain detectability for the rotating short-haired BH model. With the gradual improvement of observational techniques, future data may help refine the parameter space of this model and provide new clues for distinguishing different black hole models. This result supports the potential of the rotating short-haired BH model as a description of real universe black holes.
@@ VI Discussion and conclusions
The gravitational lensing effect has become an important tool for studying the spacetime structure of black holes. In particular, the 2019 image of M87* revealed its high consistency with the Kerr metric, strongly supporting the theoretical expectation that black holes follow the Kerr metric. However, the existence of accretion flows, dark matter halos, and non-vacuum environments poses significant challenges to the isolated Kerr black hole model. As a natural extension of the Kerr solution, the rotating short-haired BH model may provide a new theoretical framework for studying real black holes in the universe and their complex environments.
@@
This paper investigates the strong gravitational lensing effect in the rotating short-haired BH and conducts a constraint analysis on relevant parameters. We find that the hair parameter significantly influences the black hole shadow and the gravitational lensing effect. Specifically, as the intensity of the short hair increases, the event horizon radius of the rotating short-haired BH decreases, and the photon orbit radius and the impact parameter decrease with the increase of the black hole spin. Meanwhile, the photon orbit radius and the impact parameter of the Kerr BH are larger than those of the rotating short-haired BHe, and the degree of deviation increases with the increase of the hair parameter. Regarding the lensing coefficients $\bar{a}$ and $\bar{b}$, the coefficient $\bar{a}$ shows a monotonically increasing trend, while the coefficient $\bar{b}$ shows a monotonically decreasing trend, and the hair parameter exacerbates these changes.
@@
We further analyzed the observational manifestations of the rotating short-haired BH in M87* and Sgr A*. Whether for M87* or Sgr A*, the changing trends of these observables are consistent: the angular position of the relativistic image $\theta_{\infty}$ decreases as the spin parameter increases, the angular distance $s$ increases with the increase of the spin parameter, and the magnification of the relativistic image $r_{mag}$ decreases as the spin parameter increases. The presence of hair makes both $\theta_{\infty}$ and $s$ smaller than those in the case of a Kerr BH, and $r_{mag}$ is larger than that of a Kerr BH. The larger the hair parameter, the greater the degree of deviation. By analyzing the data in Table 2, it can be concluded that the hair parameter $Q_{m}$ has a remarkable impact on the angular position of the relativistic image. Its variation ranges are $22.7113\ \mu as>\theta_{\infty}(M87^{*})>13.9162\ \mu as$ and $29.3132\ \mu as>\theta_{\infty}(SgrA^{*})>17.9614\ \mu as$ respectively. These ranges are comparable to the observational results of the shadow diameters of M87* and Sgr A* obtained by the EHT Akiyama et al. (2019a, 2022b) (see Figure 7 and Table 2). Through comparing the rotating short-haired BH with the Kerr BH ($Q_{m}=0$) and the Kerr-Newman BH ($k=1$), we’ve found that the angular position deviation between the rotating short-haired BH and the Kerr BH can reach $\delta\theta_{\infty}\approx 3\ \mu as$ in the case of M87*, and $\delta\theta_{\infty}\approx 3.3786\ \mu as$ for Sgr A*. The deviations from the Kerr-Newman BH are $\delta\theta_{\infty}\approx 2.4249\ \mu as$ (for M87*) and $\delta\theta_{\infty}\approx 3.1299\ \mu as$ (for Sgr A*), respectively.
@@
In addition, the time delay effect of the rotating short-haired BH also shows significant differences from other models. When simulating M87*, the time deviation between the rotating short-haired BH and the Kerr BH can reach more than 50 hours, while the deviation from the Kerr-Newman BH can reach more than 100 hours. For Sgr A*, the time deviations are 2 minutes and 4 minutes respectively. Apparently, if the rotating short-haired BH is taken as the M87* model, the time delay effect is sufficient to distinguish it from the Kerr BH or the Kerr-Newman BH, providing theoretical support for further exploration of the properties of the hair parameter. It is worth mentioning that the maximum value of the time delay deviation does not increase monotonically with the spin parameter, but reaches its maximum at a specific spin parameter (see Figures 8 and 9 and Table 4).
@@
Finally, by integrating the observational data from the EHT, we systematically constrained the hair parameter of the rotating short-haired BH. The results indicate that, within the first confidence interval, the parameter space $(Q_{m}/M,a/M)$ of the rotating short-haired BH is highly consistent with the observational data (Figure 10). This suggests that the rotating short-haired BH cannot be excluded as a possible model for black holes in the real universe, providing theoretical support for further differentiating between different black hole models.
@@
In summary, the hair parameter has a significant impact on the strong gravitational lensing effect and the black hole shadow. These impacts may lead to certain observational differences for distinguishing between different black hole models. Future research could further explore the influence of the hair parameter on the weak gravitational lensing effect and further verify the applicability of this model through high precision observations.
@@
We acknowledge the anonymous referee for a constructive report that has significantly improved this paper. This work was supported by the Special Natural Science Fund of Guizhou University (Grant No.X2022133), the National Natural Science Foundation of China (Grant No. 12365008) and the Guizhou Provincial Basic Research Program (Natural Science) (Grant No. QianKeHeJiChu-ZK[2024]YiBan027) .

### v3

@@ V Constraints from the EHT
Figure 11: After taking the rotating short-haired BH model as a candidate model for M87* and Sgr A*, the parameter space ($Q_{m}/M,a/M$) is constrained using the EHT data. The left hand figure corresponds to M87*, and the right hand figure corresponds to Sgr A*. The red dotted line corresponds to the boundary of the $1\sigma$ confidence interval. Here, the inclination angle $\theta_{o}=\frac{\pi}{2}$.
Figure 12: Parameter space constraints for rotating short-haired black hole models as candidate models for M87* and Sgr A*. Left figure: Parameter constraints for M87*; Right figure: Parameter constraints for Sgr A*. The red dashed line represents the boundary of the corresponding $1\sigma$ confidence interval. The analysis fixes the black hole spin parameter $a=0.5$ and the observation inclination angle $\theta_{o}=\frac{\pi}{2}$.
@@
In the preceding section, we delved into the influence exerted by the hair parameter on the strong gravitational lensing phenomenon. The hair parameter acts by modifying the spacetime structure in the vicinity of the rotating short-haired BH, thereby inducing alterations in the propagation trajectories of light rays within the BH’s gravitational field. As light rays approach the event horizon, the intense gravitational field of the black hole causes the light rays to undergo deflection, giving rise to the gravitational lensing effect. This may potentially result in the formation of multiple images or regions of enhanced luminosity, as elaborated in detail in the previous section. With further deflection of the light rays, a black hole shadow is ultimately formed. The geometric properties of the shadow of a rotating short-haired BH are not only contingent upon the black hole’s spin but are also significantly influenced by the hair parameter. Consequently, while having an impact on the propagation paths of light rays, the hair parameter exerts a far-reaching influence on both the shape and size of the black hole shadow. A comprehensive discussion regarding the effect of the hair parameter on the shape of the shadow surrounding a rotating short-haired BH can be found in the reference [14]. Therefore, our emphasis does not lie in a detailed exploration of the effects on the shadow shapes. Instead, we are committed to systematically constraining the relevant parameter space of the rotating short-haired BH by leveraging the shadow data of supermassive BHs obtained from the observations of the EHT.
@@
In 2019, the EHT successfully captured the image of the black hole shadow of the supermassive BH M87* located at the center of the elliptical galaxy M87, which brought revolutionary progress to the observational and theoretical research of supermassive BHs. The EHT team measured the diameter of the black hole shadow of M87* to be $\theta_{d}=42\pm 3\ \mu\mathrm{as}$ [66, 48, 69]. This result not only provides us with precise observational evidence of the geometric structure of the BH but also offers evidence for verifying the validity of general relativity in extreme gravitational fields. Subsequently, in 2022, the EHT team further enhanced the observational technology and data processing capabilities and successfully measured the diameter of the black hole shadow of the supermassive BH Sgr A* located at the center of the Milky Way. The result is $\theta_{d}=51.8\pm 2.3\ \mu\mathrm{as}$ [49]. These data play a crucial role in the detection of gravitational theories and parameter constraints, such as [70, 71, 72, 73, 74, 71].
@@
For axisymmetric black holes, the geometric shape of the black hole shadow is influenced by the spin parameter. To more accurately describe the shape of the shadow, an astronomical coordinate system [75] is usually introduced, which is defined as
@@
Here, the prime (’) indicates the derivative with respect to $r$, and $\Delta$ is a function term associated with the rotating short-haired BH. The detailed derivation process of this part can be found in corresponding literature, such as [14, 71, 76], etc.
Once we obtain the corresponding expressions for celestial coordinates, it becomes straightforward to determine the shadow shape of a rotating short-haired black hole. This aspect has been extensively discussed in the literature [14]; therefore, we shall not elaborate further on this topic here. The primary objective of this section is to constrain the hair parameter and spin parameter of the rotating short-haired black hole using the supermassive black hole shadow data observed by the EHT.
If the rotating short-haired BH model is considered as a candidate for M87* or SgrA*, then for a distant observer, the angular diameter $\theta_{d}$ of the shadow image around the rotating short-haired BH can be expressed as (e.g., [77, 71, 72, 78], etc.)
$\theta_{d}=\frac{2R_{a}}{d},\quad R_{a}^{2}=\frac{A}{\pi}.$ (53)
Where $d$ represents the distance to the observer, $R_{a}$ is the area shadow radius, and $A$ is the region within the shadow contour [76, 79]
[…]
where $r_{+}$ and $r_{-}$ are the roots of $Y^{2}=0$ outside the event horizon radius, corresponding to the prograde and retrograde radii of stable circular orbits.
@@
Furthermore, in order to explore the impact of parameter $k$ on the constraint of the hair parameter $Q_{m}$, we fixed the spin parameter $a=0.5$ in Figure 12 and systematically analyzed the changes in the constraint interval of $Q_{m}$ under different values of $k$. The results show that, regardless of whether it is based on the observational data constraints from M87* or Sgr A*, an increase in parameter $k$ leads to a remarkable weakening of the constraining ability on the hair parameter $Q_{m}$. (The red dashed line in the figure represents the $1\sigma$ confidence interval boundary.) Specifically, as the value of $k$ increases, the allowable interval for $Q_{m}$ gradually broadens, and the constraining power gradually diminishes. This conclusion is in high agreement with the analysis results regarding the influence of $k$ on observables such as the light deflection angle and the relativistic angular position of the image in the previous sections. It further validates that a degeneracy exists between the rotating short-haired BH and the Kerr BH in the region of large $k$ values.
Overall, whether using the shadow data of M87* or Sgr A* to constrain the parameter space $(Q_{m}/M,a/M)$ of the rotating short-haired BH model, our analysis results show that the parameter space of the rotating short-haired BH model is highly consistent with the range of the observational data from the EHT. Therefore, through a systematic analysis of the EHT observational results of M87* and Sgr A* within the first confidence interval, we cannot rule out the rotating short-haired BH model as a potential candidate for real universe BHs. In addition, although the current observational accuracy still has limitations, the EHT data has provided a certain detectability for the rotating short-haired BH model. With the gradual improvement of observational techniques, future data may help refine the parameter space of this model and provide new clues for distinguishing different BH models. This result supports the potential of the rotating short-haired BH model as a description of real universe BHs.
@@ VI Discussion and conclusions
The gravitational lensing effect has become an important tool for studying the spacetime structure of BHs. In particular, the 2019 image of M87* revealed its high consistency with the Kerr metric, strongly supporting the theoretical expectation that black holes follow the Kerr metric. However, the existence of accretion flows, dark matter halos, and non-vacuum environments poses significant challenges to the isolated Kerr BH model. As a natural extension of the Kerr solution, the rotating short-haired BH model may provide a new theoretical framework for studying real BHs in the universe and their complex environments.
@@
This paper investigates the strong gravitational lensing effect in the rotating short-haired BH and conducts a constraint analysis on relevant parameters. We find that the hair parameter significantly influences the black hole shadow and the gravitational lensing effect. Specifically, as the intensity of the short hair increases, the event horizon radius of the rotating short-haired BH decreases, and the photon orbit radius and the impact parameter decrease with the increase of the black hole spin. Meanwhile, the photon orbit radius and the impact parameter of the Kerr BH are larger than those of the rotating short-haired BHe, and the degree of deviation increases with the increase of the hair parameter. For the lensing coefficients $\bar{a}$ and $\bar{b}$, they exhibit monotonically increasing and decreasing trends respectively. The hair parameter intensifies this variation, while parameter $k$ suppresses the influence effect of the hair parameter.
@@
We further analyzed the observational manifestations of the rotating short-haired BH in M87* and Sgr A*. Whether for M87* or Sgr A*, the changing trends of these observables are consistent: the angular position of the relativistic image $\theta_{\infty}$ decreases as the spin parameter increases, the angular separation $S$ increases with the increase of the spin parameter, and the magnification of the relativistic image $r_{mag}$ decreases as the spin parameter increases. The existence of the hair parameter $Q_{m}$ causes both $\theta_{\infty}$ and $S$ to be smaller than in the Kerr BH case, while $r_{mag}$ is larger than in the Kerr BH case. As the hair parameter $Q_{m}$ increases, this deviation gradually strengthens; however, an increase in parameter $k$ suppresses this effect, causing these observables to gradually approach the Kerr BH case. By analyzing the data in Table 2, it can be concluded that the hair parameter $Q_{m}$ has a remarkable impact on the angular position of the relativistic image. Its variation ranges are $22.7113\ \mu as>\theta_{\infty}(M87^{*})>13.9162\ \mu as$ and $29.3132\ \mu as>\theta_{\infty}(SgrA^{*})>17.9614\ \mu as$ respectively. These ranges are comparable to the observational results of the shadow diameters of M87* and Sgr A* obtained by the EHT [48, 80] (see Figure 7 and Table 2). Through comparing the rotating short-haired BH with the Kerr BH ($Q_{m}=0$) and the Kerr-Newman BH ($k=1$), we’ve found that the angular position deviation between the rotating short-haired BH and the Kerr BH can reach $\delta\theta_{\infty}\approx 3\ \mu as$ in the case of M87*, and $\delta\theta_{\infty}\approx 3.3786\ \mu as$ for Sgr A*. The deviations from the Kerr-Newman BH are $\delta\theta_{\infty}\approx 2.4249\ \mu as$ (for M87*) and $\delta\theta_{\infty}\approx 3.1299\ \mu as$ (for Sgr A*), respectively.
@@
In addition, the time delay effect of the rotating short-haired BH also shows significant differences from other models. When simulating M87*, the time deviation between the rotating short-haired BH and the Kerr BH can reach more than $50$ hours, while the deviation from the Kerr-Newman BH can reach more than $100$ hours. For Sgr A*, the time deviations are $2$ minutes and $4$ minutes respectively. Apparently, if the rotating short-haired BH is taken as the M87* model, the time delay effect is sufficient to distinguish it from the Kerr BH or the Kerr-Newman BH, providing theoretical support for further exploration of the properties of the hair parameter. It is worth mentioning that the maximum value of the time delay deviation does not increase monotonically with the spin parameter, but reaches its maximum at a specific spin parameter (see Figures 9 and 10 and Table 5).
@@
Finally, by integrating the observational data from the EHT, we systematically constrained the hair parameter of the rotating short-haired BH. The results indicate that, within the first confidence interval, the parameter space $(Q_{m}/M,a/M)$ of the rotating short-haired BH is highly consistent with the observational data (Figure 11). This suggests that the rotating short-haired BH cannot be excluded as a possible model for BHs in the real universe, providing theoretical support for further differentiating between different BH models.
@@
In summary, the hair parameter has a significant impact on the strong gravitational lensing effect and the black hole shadow. These impacts may lead to certain observational differences for distinguishing between different BH models. Future research could further explore the influence of the hair parameter on the weak gravitational lensing effect and further verify the applicability of this model through high precision observations.
@@
We acknowledge the anonymous referee for a constructive report that has significantly improved this paper.This work was supported by Guizhou Provincial Basic Research Program(Natural Science)(Grant No.QianKeHeJiChu-[2024]Young166), the Special Natural Science Fund of Guizhou University (Grant No.X2022133), the National Natural Science Foundation of China (Grant No.12365008) and the Guizhou Provincial Basic Research Program (Natural Science) (Grant No.QianKeHeJiChu-ZK[2024]YiBan027 and QianKeHeJiChu-MS[2025]680).